OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (x-5)^2/144 - (y-4)^2/81 = 1.

OpenStudy (anonymous):

Vertices: (14, 4), (-4, 4); Foci: (-4, 4), (14, 4) Vertices: (17, 4), (-7, 4); Foci: (-10, 4), (20, 4) Vertices: (4, 17), (4, -7); Foci: (4, -10), (4, 20) Vertices: (4, 14), (4, -4); Foci: (4, -4), (4, 14)

OpenStudy (e.mccormick):

Do you know how to find any of those?

OpenStudy (anonymous):

no i would like to know how to solve the problem

OpenStudy (e.mccormick):

OK. The first thing to note is what is being subtracted includes a y. That means the hyperbola is basically this shape: 

OpenStudy (e.mccormick):

The standard form is: \(\dfrac{(x-h)^2}{a^2}-\dfrac{(y-k)^2}{b^2}=1\) Now, there is also a c involved. \(a^2+b^2=c^2\)  The b values are used to find the asymptotes.